Bargain Rental Car offers rental cars in an off-airport location near a major tourist destination in California. Management would like to better understand the variable and fixed portions of it car washing costs. The company operates its own car wash facility in which each rental car that is returned is thoroughly cleaned before being released for rental to another customer. Management believes that the variable portion of its car washing costs relates to the number of rental returns. Accordingly, the following data have been compiled: Month Rental Returns Car Wash Costs January 2,400 $ 10,900 February 2,500 $ 13,100 March 2,700 $ 11,700 April 3,000 $ 14,100 May 3,600 $ 16,100 June 5,000 $ 23,100 July 5,500 $ 22,100 August 5,500 $ 21,900 September 4,700 $ 22,700 October 4,000 $ 20,800 November 2,200 $ 10,600 December 2,800 $ 13,900 2. Using least-squares regression, estimate the variable cost per rental return and the monthly fixed cost incurred to wash cars. (Round Fixed cost to the nearest whole dollar amount and the Variable cost per unit to 2 decimal places.)
Answer #1

Ans: The solution is as follows :-

Firstly arrange the data in a tabular format -

Month No. of cars Total Cost
January 2400 10900
February 2500 13100
March 2700 11700
April 3000 14100
May 3600 16100
June 5000 23100
July 5500 22100
August 5500 21900
September 4700 22700
October 4000 20800
November 2200 10600
December 2800 13900

The formulas used here are as follows, and all the components will be worked out from the table made above.

n (Σxy) - (Σx)(Σy) Variable cost per unit = b =- n (Σx?)-(Σx)? (Σy)-b(Σx) Fixed cost = a=-- n

Table showing calculation of Variable cost per unit -

S.no. x y x^2 xy
1 2400 10900 5760000 26160000
2 2500 13100 6250000 32750000
3 2700 11700 7290000 31590000
4 3000 14100 9000000 42300000
5 3600 16100 12960000 57960000
6 5000 23100 25000000 115500000
7 5500 22100 30250000 121550000
8 5500 21900 30250000 120450000
9 4700 22700 22090000 106690000
10 4000 20800 16000000 83200000
11 2200 10600 4840000 23320000
12 2800 13900 7840000 38920000
Total 43900 201000 177530000 800390000

Here,

n = 12

∑ x = number of cars = 43900

∑ y = Total cost = 201000

∑ x^2 = 177530000

∑ xy = 800390000

So after applying the given values in formula we get,

Variable cost per unit = 12*( 800390000) - (43900)*(201000)
12*(117530000) - (43900)^2
=(9604680000 - 8823900000 )
(2130360000 - 1927210000)
= (780780000)
(203150000)
= $ 3.84 / car

Now calculating Fixed cost -

Here,
∑ y = 201000
b = Variable cost per unit = $ 3.84
∑ x = 43900
n = 12
Applying the above values in formula we get,

Fixed cost = ( 201000) - (3.84*43900)
12
= (32424) / 12 = $ 2702